OpenStudy (anonymous):

Find the local maximum and minimum values of f using both the First and Second Derivative Tests. f(x) = 4 + 12x^2 − 8x^3

OpenStudy (anonymous):

\[f(x)=4+12x ^{2}-8x ^{3}\] \[f'(x)=24x-24x ^{2}=24x \left( 1-x \right)\] f'(x)=0 gives x=0,1 f''(x)=24-48x atx=0,f''(x)=24>0,there is local minima at x=0 and minimum value of f(x)=4 at x=1,f''(x)=-24<0,there is local maxima at x=1 and maximum value=4+12-8=8

OpenStudy (anonymous):

first derivative test when x<0 slightly f'(x)=-(+)=-ve when x>0 slightly f'(x)=+(+)=+ve f'(x) changes sign from -ve to +ve as x passes through 0,hence there is local minima at x=0 when x<1 slightly,x-1<0 or1-x>0 f'(x)=+(+)=+ve when x>1 slightly,x-1>0 or 1-x<0 f'(x)=+(-)=-ve f'(x) changes sign from +ve to-ve as x passes through 1,hence there is local maxima at x=1